Solve Problems Using Linear Systems
13. A plane took 4h to fly 2200 km from Saskatoon to Toronto with a tail wind. The return trip, with a head wind, took 5h. Find the speed of the plane in still air and the wind speed.
Can you please help me with this word problem?
thanksss

Answer by checkley77(12844)

D=RT or D/T=R
2200/4=R
550=R IS THE SPEED WITH THE WIND.
2200/5=T
440=R IS THE SPEED AGAINST THE WIND.
(550-440)/2=110/2=55 MPH IS THE WIND SPEED.
550-55=495 MPH IS THE PLANE SPEED
440+55=495 DITTO.